OpenStudy (anonymous):

Nitroglycerin detonates according to the balanced reaction: 4 C3H5N3O9(l) → 12 CO2(g) + 6 N2(g) + O2(g) + 10 H2O(g). If 150.0 grams of nitroglycerin is detonated, how many L of products (assume all gases at standard temperature and pressure, so 1.00 mole = 22.4 L of volume) are released?

OpenStudy (anonymous):

Help please...I dont know how to approch this problem.

OpenStudy (jfraser):

you know that 4mol of nitroglycerin will produce a total of 29 moles of mixed products, the balanced reaction says so. Find how many moles of reactant you have, then use that mole ratio to find the total moles of products formed. Use the molar volume as the last conversion factor to get total volume of the products.\[150.0g Nitro * (\frac{1mol nitro}{227g nitro})*(\frac{29mol products}{4mol nitro})*(\frac{22.4L}{1mol})\]

OpenStudy (anonymous):

Life makes sense now. LOL Thanks so much. Basically the 22.4 L is all of the products!! I didnt catch that at first.